Can you please help me find the slope of hte equation 6x-5y=30.
Answer by tutorcecilia(2105)

6x-5y=30 [Standard form of a line]
y=mx+b [slope-intercept form of a line]
-5y=-6x+30 [re-write in the slope-intercept form of a line]
y=6/5x-6
y=any y value on the line
m=the slope of the line
x=any corresponding x value on the line
b = the y-intercept of the line
.
in this case, m = 6/5